# Filtrations on derived category of twisted K3 surfaces

### v1

@@ Acknowledgement
We want to thank Ziyu Zhang and Junliang Shen for helpful discussions. The first author is supported by NSFC grant (No. 12121001, No. 12171090) and Shanghai Pilot Program for Basic Research (No. 21TQ00).
@@
In this paper, we use $\operatorname{CH}^{\bullet}(X)$ to denote the Chow groups of $X$ with rational coefficients.
@@ 2.1. Twisted K3 surface and twisted sheaves
A $d$-fold $\mathscr{X}$-twisted sheaf on $p:\mathscr{X}\to X$ is an $\mathcal{O}_{\mathscr{X}}$-module of weight $d$ with respect to the $\mu_{m}$-action (cf[16, Def 2.1.2.2]). Let $\mathrm{D}^{(d)}(\mathscr{X})$ be the bounded derived category of coherent $d$-fold $\mathscr{X}$-twisted sheaves. For a $\mathscr{X}$-twisted locally free sheaf $\mathcal{E}\in\mathrm{D}^{(1)}(\mathscr{X})$, we have a diagram
[…]
where $\mathbb{P}(\mathcal{E})\to Y$ is essentially trivial and $Y$ is served as a Severi-Brauer variety of $(X,\alpha)$.
@@ 2.2. Chern character of twisted sheaves
Let $K_{0}(\mathscr{X})$ be the Grothendieck group of $1$-fold $\mathscr{X}$-twisted coherent sheaves. In [17, §3.3.4], Lieblich-Maulik-Snowden have introduced a twisted Chern character
@@
accordingly. The definition above avoids the use of the Chow rings of stacks. However, one can compare it with the Chern character $\operatorname{ch}:K_{0}(\mathscr{X})\to\operatorname{CH}^{\ast}(\mathscr{X})$ for stacks. An easy fact is
@@
[…]
@@
In particular, we have $c_{2}(\Phi_{\ast}\mathcal{L})=[C]^{2}-\phi_{\ast}(c_{1}(\mathcal{L}))\in\operatorname{CH}_{0}(X)$. It has the same form as in the untwisted case. The foregoing argument uses the Grothendieck-Riemann-Roch theorem on $\mu_{m}$-gerbes (cf[16, Proposition 2.2.7.8]). Alternatively one can obtain the above formula using the same method as [4, Lemma 4.1.4].
@@
such that $[\mathscr{X}]=\delta(B)$, where $\delta:\mathrm{H}^{2}(X,\mathbb{Q})\to\mathrm{H}^{2}(X,\mathcal{O}_{X}^{\times})$ is induced from the exponential exact sequence. We define
@@
Up to a lattice isometry, $\widetilde{\mathrm{H}}(\mathscr{X})$ is independent of the choice of a $B$-field lift. One can similarly define the twisted algebraic Mukai lattice $\widetilde{\mathrm{H}}_{\rm alg}(\mathscr{X})$ by replacing $\widetilde{\mathrm{H}}(X)$ with $\widetilde{\mathrm{H}}_{\rm alg}(X)$. For $\mathcal{F}\in\mathrm{D}^{b}(\mathscr{X})$, its Mukai vector is defined by
@@ 2.5. Bridgeland moduli space
Let $\mathbf{v}=(r,L,s)\in\widetilde{\mathrm{H}}(\mathscr{X})$ be a primitive Mukai vector with $\mathbf{v}^{2}\geq 0$ and let $\sigma$ be a $\mathbf{v}$-generic Bridgeland stability condition. Let $\mathscr{M}_{\sigma}(\mathscr{X},\mathbf{v})$ be the moduli stack of $\sigma$-stable objects in $\mathrm{D}^{b}(\mathscr{X})$ with Mukai vector $\mathbf{v}$. Let $\mathscr{M}_{\sigma}(\mathscr{X},\mathbf{v})\to M_{\sigma}(\mathscr{X},\mathbf{v})$ be the coarse moduli map, which can be also viewed as a gerbe on $M_{\sigma}(\mathscr{X},\mathbf{v})$. In this way, we have a universal object
@@
According to [27, Sections 7 and 8], [5, Proposition 14.2] and [2, Theorem 6.10, Section 7], when $\mathbf{v}^{2}\geq-2$, $M_{\sigma}(\mathscr{X},\mathbf{v})$ is a non-empty smooth hyper-Kähler variety of K3${}^{[n]}$-type of dimension $2+v^{2}$. If $\sigma$ is a Gieseker stability condition with respect to an ample line bundle $H\in\operatorname{Pic}(X)$, we may write it as $\mathscr{M}_{H}(\mathscr{X},\mathbf{v})$. When $r\geq 0$ and $\mathbf{v}^{2}\geq-2$, the moduli space $\mathscr{M}_{H}(\mathscr{X},\mathbf{v})$ is not empty for a $\mathbf{v}$-generic $H$. There is a natural isometry
@@
induced by the (quasi-)universal object, where $\mathbf{v}^{\perp}$ is the orthogonal complement of $\mathbf{v}$ in $\widetilde{\mathrm{H}}_{\rm alg}(\mathscr{X})$ (cf[27, Sections 7 and 8]).
@@
Theorem 2.7.
[2, Theorem 1.1, Theorem 1.5] With notations as above. Assume that there exists an isotropic class $\mathbf{w}\in\mathbf{v}^{\perp}$, then $M_{\sigma}(\mathscr{X},\mathbf{v})$ admits a birational Lagrangian fibration.
3. Cycles on HK with birational Lagrangian fibration
@@ 3.1. Density of constant cycle Lagrangian subvarieties
Let us recall some general results for cycles on hyper-Kähler varieties with a Lagrangian fibration. The first result is a strengthen of [18, Theorem 1.3]
@@
As the property is preserved under birational transformations, we are reduced to consider the case $M$ admits a Lagrangian fibration $M\to\mathbb{P}^{n}$. If $T\to\mathbb{P}^{n}$ is a morphism, we set
@@
Let $U\subseteq\mathbb{P}^{n}$ be the open locus parametrizing all the smooth fibers of $M\to\mathbb{P}^{n}$. According to [18, Theorem 1.3], there is a constant cycle Lagrangian subvariety $\Sigma\subseteq M$ such that the restriction $\Sigma_{U}=\Sigma\cap M_{U}\to U$ is finite. This gives a map
@@ 3.4. Twisted torsion sheaves with maximal dimensional rational equivalent orbit
By Example 2.4, $c_{2}(\mathcal{E})$ is a zero cycle supported on a curve $C\in|L|$. Since the genus of $C$ is less than or equal to $\frac{1}{2}(L^{2}+2)=\mathrm{d}(\mathcal{E})$, the assertion then follows from [20, Claim 0.2] (see also [26, Proposition 15]). ∎
@@
Due to Proposition 3.5, $\pi_{1}$ is dominant. By the same argument in [20, Proposition 1.3], we know that $\pi_{2}$ is also dominant.
@@
Then for any $\mathcal{F}\in\pi_{2}(\pi_{1}^{-1}(\Sigma_{\mathscr{X}}^{\prime}))$, we have $c_{2}(\mathcal{F})\in\mathbf{S}_{0}(X)$ by [26, Theorem 9]. Therefore $c_{2}(\mathcal{E}^{\prime})\in\mathbf{S}_{0}(X)$ by the definition of $\Gamma$ for any $\mathcal{E}^{\prime}\in\Sigma_{\mathscr{X}}^{\prime}$. ∎
@@ 3.7. Correspondence between torsion sheaves
If $\mathscr{X}\to X$ is essentially trivial, this is established in [15, Theorem 3.1 & Theorem 1.1]. If $\mathscr{X}\to X$ is not essentially trivial, we can establish a finite correspondence between $M_{X}^{\operatorname{sm}}$ and $M_{\mathscr{X}}^{\operatorname{sm}}$ which preserves the two filtrations. More precisely, consider the incidence variety
@@
with $\dim Z\geq\mathrm{d}(\mathcal{E})-i$. Then $Z^{\prime}=\pi_{2}(\pi_{1}^{-1}(Z))$ is a constant cycle subvariety by Theorem 2.6 and we have $\dim Z^{\prime}=\dim Z\geq\frac{\mathbf{v}^{2}+2}{2}-i$. This means $\dim O_{E}\geq\frac{\mathbf{v}^{2}+2}{2}-i$ for any $E\in Z^{\prime}$. According to the proof of [15, Theorem 1.1], we have
@@
Let us explain the construction of $W$. Let $\Sigma_{\mathscr{X}}\subseteq M_{\mathscr{X}}$ and $\Sigma_{X}\subseteq M_{X}$ be a constant cycle Lagrangian subvariety defined in the proof of Proposition 3.2. Their restrictions to the smooth locus $M_{\mathscr{X}}^{\operatorname{sm}}$ (resp. $M_{X}^{\operatorname{sm}}$) are multisections of the abelian fibration of degree $d_{1}$ (resp. $d_{2}$).
@@ 4.1. Construction of $\mathfrak{o}_{\mathscr{X}}$
$\mathfrak{o}_{\mathscr{X}}=\frac{c_{2}(\mathcal{F})}{\operatorname{rank}\mathcal{F}}+(1-\frac{\deg(\mathcal{F})}{\operatorname{rank}\mathcal{F}})\mathfrak{o}_{X}.$ (18)
to be a zero cycle of degree $1$.
@@
Let $\mathbf{v}\in\widetilde{\mathrm{H}}(\mathscr{X})$ and $\mathbf{v}^{\perp}\subseteq\widetilde{\mathrm{H}}_{\rm alg}(\mathscr{X})$ contains an isotropic element. Assume that $\operatorname{rank}(\mathbf{v})=r>\frac{\mathbf{v}^{2}+2}{2}$ and there exists a slope stable element in $M_{H}(\mathscr{X},\mathbf{v})$. Then $M_{H}(\mathscr{X},\mathbf{v})$ contains a constant cycle Lagrangian subvariety $Z$ such that $Z\cap M_{H}^{\rm l.f.}(\mathscr{X},\mathbf{v})\neq\emptyset$.
@@
Note that under the condition $r>\frac{\mathbf{v}^{2}+2}{2}$, every $\mu_{H}$-stable $\mathscr{X}$-twisted sheaves are locally free (cf[27, Remark 3.2] and [11, Remark 6.1.9]) and the assertion follows from our assumption.
@@
The existence of slope stable elements can be easily achieved. Let $r_{0}$ be the minimal rank of $\mathscr{X}$-twisted sheaves in $\mathrm{Coh}^{(1)}(\mathscr{X})$. Then $r_{0}$ is the order of $[\mathscr{X}]$ in $\operatorname{Br}(X)$ and every element in $\mathrm{Coh}^{(1)}(\mathscr{X})$ has rank $ar_{0}$ for some $a\in\mathbb{Z}$ (cf. [28, Lemma 3.2]). For any primitive vector
$\mathbf{v}=(ar_{0},h,s)\in\widetilde{\mathrm{H}}(\mathscr{X})$
with $a>0$ and $\gcd(a,H\cdot h)=1$, we have $\mathcal{E}\in M_{H}(\mathscr{X},\mathbf{v})$ is $\mu_{H}$-stable. When $r>\frac{\mathbf{v}^{2}+2}{2}$, we actually have $M_{H}^{\rm l.f.}(\mathscr{X},\mathbf{v})=M_{H}(\mathscr{X},\mathbf{v})$.
Moreover, to see the existence of $\mathbf{v}$ whose orthogonal complement contains an isotropic element, one can just let $s=0$ and the isotropic vector $(1,0,0)$ is contained in $\mathbf{v}^{\perp}$.
Now we show that $\mathfrak{o}_{\mathscr{X}}$ is independent of the choice of the moduli space $N$ and the constant cycle Lagrangian subvariety $Z$. The key result is
@@
Remark 4.8 (Descent property).
We speculate that $\mathfrak{o}_{\mathscr{X}}=\mathfrak{o}_{X}$. The main evidence is that due to our construction, the twisted BV class $\mathfrak{o}_{\mathscr{X}}$ has nice descent property. This follows from the descent property of constant cycle Lagrangian subvarieties (cf. [8, Proposition 3.6]).
Here we can give a more explicit proof. Suppose $X$ is defined over $\bar{k}$. Take a Mukai vector $\mathbf{v}=(r,h,s)$ such that $\mathbf{v}^{2}=0$ and $r>1$. Then $M=M_{H}(\mathscr{X},\mathbf{v})$ is a K3 surface defined over $\bar{k}$ and it consists of Gieseker stable $\mathscr{X}$-twisted locally free sheaves. Take a $\bar{k}$-rational curve $C\subseteq M$ and let $\mathcal{F}$ be a point on $C$. Then $\mathfrak{o}_{\mathscr{X}}=\frac{c_{2}(\mathcal{F})}{r}+(1-\frac{\deg(\mathcal{F})}{r})\mathfrak{o}_{X}$ is clearly defined over $\bar{k}$.
When $\bar{k}=\bar{\mathbb{Q}}$, the Bloch-Beilinson conjecture predicts $\operatorname{CH}_{0}(X_{\bar{\mathbb{Q}}})=\mathbb{Z}\mathfrak{o}_{X}$ and $\mathfrak{o}_{\mathscr{X}}=\mathfrak{o}_{X}$.
@@
The idea is to consider the degenerate loci of morphisms between twisted sheaves. Let us recall the definition and some basic properties.
@@
Let $\mathcal{E}$ and $\mathcal{F}$ be two $\mathscr{X}$-twisted locally free sheaves. For $t:\mathcal{E}\to\mathcal{F}\in\operatorname{Hom}(\mathcal{E},\mathcal{F})$, we define the $k$-th degenerate locus $\mathscr{D}_{k}(t)$ to be the closed substack of $\mathscr{X}$ defined by the image of the map
@@
We can also define the schematic $k$-th degenerate locus $D_{k}(t)$ which is the scheme-theoretic image in $X$. (cf[16, Definition 2.2.6.2])
Form the definition, since $\mathscr{H}om(\wedge^{k+1}\mathcal{E},\wedge^{k+1}\mathcal{F})$ is the pullback of a sheaf on $X$, $\mathscr{D}_{k}(t)\to D_{k}(t)$ is the pullback of $\mathscr{X}\to X$ via $D_{k}(t)\to X$. The following result is well-known for untwisted locally free sheaves (cf[6, §4.1][12, Lemma 2.14]).

### v3

@@ Acknowledgement
We want to thank Ziyu Zhang and Junliang Shen for helpful discussions. Z. Chen and R. Zhang are supported by the NKRD Program of China (No. 2020YFA0713200) and NSFC grant (No. 12121001). Z. Li is supported by NSFC grant (No. 12171090 and No. 12425105) and Shanghai Pilot Program for Basic Research (No. 21TQ00). Z. Li is also a member of LMNS.
@@
In this paper, we use $\operatorname{CH}^{\bullet}(X)$ or $\operatorname{CH}_{\bullet}(X)$ to denote the Chow groups of $X$ with rational coefficients.
@@ 2.1. Twisted K3 surface and twisted sheaves
A $d$-fold $\mathscr{X}$-twisted sheaf on $p:\mathscr{X}\to X$ is an $\mathcal{O}_{\mathscr{X}}$-module of weight $d$ with respect to the $\mu_{m}$-action (cf[18, Def 2.1.2.2]). Let $\mathrm{D}^{(d)}(\mathscr{X})$ be the bounded derived category of coherent $d$-fold $\mathscr{X}$-twisted sheaves. For a $\mathscr{X}$-twisted locally free sheaf $\mathcal{E}\in\mathrm{D}^{(1)}(\mathscr{X})$, we have a diagram
${\mathbb{P}(\mathcal{E})}$${Y}$${\mathscr{X}}$${X,}$$\scriptstyle{q}$$\scriptstyle{\varpi}$$\scriptstyle{\pi}$$\scriptstyle{p}$
where $\mathbb{P}(\mathcal{E})\to Y$ is essentially trivial and $Y$ is served as a Severi-Brauer variety of $(\mathscr{X})$.
@@ 2.2. Chern character of twisted sheaves
Let $K_{0}(\mathscr{X})$ be the Grothendieck group of $1$-fold $\mathscr{X}$-twisted coherent sheaves. In [19, §3.3.4], Lieblich-Maulik-Snowden have introduced a twisted Chern character
@@
accordingly. The definition above avoids the use of the Chow groups of stacks. However, one can compare it with the Chern character $\operatorname{ch}:K_{0}(\mathscr{X})\to\operatorname{CH}^{\ast}(\mathscr{X})$ for stacks. An easy fact is
@@
${\mathscr{C}}$${\mathscr{X}}$${C}$${X.}$$\scriptstyle{\Phi}$$\scriptstyle{q}$$\scriptstyle{p}$$\scriptstyle{\phi}$
@@
In particular, we have $c_{2}(\Phi_{\ast}\mathcal{L})=[C]^{2}-\phi_{\ast}(c_{1}(\mathcal{L}))\in\operatorname{CH}_{0}(X)$. It has the same form as in the untwisted case. The foregoing argument uses the Grothendieck-Riemann-Roch theorem on $\mu_{m}$-gerbes (cf[18, Proposition 2.2.7.8]). Alternatively one can obtain the above formula using the same method as [5, Lemma 4.1.4].
@@
such that $[\mathscr{X}]=\delta(\mathbf{B})$, where $\delta:\mathrm{H}^{2}(X,\mathbb{Q})\to\mathrm{H}^{2}(X,\mathcal{O}_{X}^{\times})$ is induced from the exponential map. We define
@@
Up to a lattice isometry, $\widetilde{\mathrm{H}}(\mathscr{X})$ is independent of the choice of a $B$-field lift. One can similarly define the twisted algebraic Mukai lattice $\widetilde{\mathrm{H}}_{\rm alg}(\mathscr{X})$ by replacing $\widetilde{\mathrm{H}}(X)$ with $\widetilde{\mathrm{H}}_{\rm alg}(X)$. For $\mathcal{F}\in\mathrm{D}^{(1)}(\mathscr{X})$, its Mukai vector is defined by
@@ 2.4. Bridgeland moduli space
Let us fix some notations. Recall that a Bridgeland stability condition $\sigma$ on $\mathrm{D}^{(1)}(\mathscr{X})$ consists of a pair $(Z,A)$, where $Z\colon K_{\mathrm{num}}(\mathscr{X})\to\mathbb{C}$ is a central charge and $A\subset\mathrm{D}^{(1)}(\mathscr{X})$ is the heart of a bounded $t$-structure, satisfying certain properties (see [3, Section 2]).
For $\phi\in\mathbb{R}$, denote by $P(\phi)$ the full subcategory of objects $\mathcal{E}$ with $Z(\mathcal{E})\in\mathbb{R}_{>0}\cdot e^{\pi i\phi}$. Then $P$ is called a slicing of $\mathrm{D}^{(1)}(\mathscr{X})$, and $\phi$ is called the phase of $\mathcal{E}$. The stability condition $\sigma$ is called geometric if there exist $\omega,\beta\in\mathrm{NS}(\mathscr{X})\otimes\mathbb{Q}$ with $\omega$ ample, such that
$Z=Z_{\omega,\beta}\colon K_{\mathrm{num}}(\mathscr{X})$ $\longrightarrow\mathbb{C},$
$\mathcal{F}$ $\mapsto\langle\exp(\beta+i\omega),\mathbf{v}(\mathcal{F})\rangle,$
and the heart $A_{\omega,\beta}=\langle F_{\omega,\beta},T_{\omega,\beta}[1]\rangle$ is obtained by tilting with respect to the torsion pair $(T_{\omega,\beta},F_{\omega,\beta})$, where
[…]
Here, $\mu_{\omega,\beta}(\mathcal{F})=\frac{c_{1}(\mathcal{F})\cdot\omega}{\operatorname{rank}(\mathcal{F})}$ denotes the twisted slope. The set of Bridgeland stability conditions is denoted by $\mathrm{Stab}(\mathscr{X})$, equipped with a natural metric topology. There is a distinguished connected component containing all geometric stability conditions ([2, Theorem 2.10]), denoted by $\mathrm{Stab}^{\dagger}(\mathscr{X})$.
Let $\mathbf{v}=(r,L,s)\in\widetilde{\mathrm{H}}(\mathscr{X})$ be a primitive Mukai vector with $\mathbf{v}^{2}\geq 0$ and let $\sigma$ be a $\mathbf{v}$-generic Bridgeland stability condition. Let $\mathscr{M}_{\sigma}(\mathscr{X},\mathbf{v})$ be the moduli stack of $\sigma$-stable objects in $\mathrm{D}^{(1)}(\mathscr{X})$ with Mukai vector $\mathbf{v}$. Let $\mathscr{M}_{\sigma}(\mathscr{X},\mathbf{v})\to M_{\sigma}(\mathscr{X},\mathbf{v})$ be the coarse moduli map, which can be also viewed as a gerbe on $M_{\sigma}(\mathscr{X},\mathbf{v})$. In this way, we have a universal object
@@
According to [29, Sections 7 and 8], [6, Proposition 14.2] and [2, Theorem 6.10, Section 7], when $\mathbf{v}^{2}\geq-2$, $M_{\sigma}(\mathscr{X},\mathbf{v})$ is a non-empty smooth hyper-Kähler variety of K3[n]-type of dimension $2+v^{2}$. If $\sigma$ is a Gieseker stability condition with respect to an ample line bundle $H\in\operatorname{Pic}(X)$, we may write it as $\mathscr{M}_{H}(\mathscr{X},\mathbf{v})$. When $r\geq 0$ and $\mathbf{v}^{2}\geq-2$, the moduli space $\mathscr{M}_{H}(\mathscr{X},\mathbf{v})$ is not empty for a $\mathbf{v}$-generic $H$. There is a natural isometry
@@
induced by the (quasi-)universal object, where $\mathbf{v}^{\perp}$ is the orthogonal complement of $\mathbf{v}$ in $\widetilde{\mathrm{H}}_{\rm alg}(\mathscr{X})$ (cf[29, Sections 7 and 8]).
@@
Theorem 2.4 ([2, Theorem 1.1, Theorem 1.5]).
Notations as above. Assume that there exists an isotropic class $\mathbf{w}\in\mathbf{v}^{\perp}$, then $M_{\sigma}(\mathscr{X},\mathbf{v})$ admits a birational Lagrangian fibration.
3. Cycles on HK with birational Lagrangian fibrations
@@ 3.1. Density of constant cycle Lagrangian subvarieties
Let us recall some general results for cycles on hyper-Kähler varieties with a Lagrangian fibration. The first result is a strengthening of [20, Theorem 1.3] .
@@
As the property is preserved under birational transformations, we are reduced to considering the case that $M$ admits a Lagrangian fibration $M\to\mathbb{P}^{n}$. If $T\to\mathbb{P}^{n}$ is a morphism, we set
@@
Let $U\subseteq\mathbb{P}^{n}$ be the open locus parametrizing all the smooth fibers of $M\to\mathbb{P}^{n}$. According to [20, Theorem 1.3], there is a constant cycle Lagrangian subvariety $\Sigma\subseteq M$ such that the restriction $\Sigma_{U}=\Sigma\cap M_{U}\to U$ is finite. This gives a map
@@ 3.2. Twisted torsion sheaves with maximal dimensional rational equivalent orbit
By Example 2.3, $c_{2}(\mathcal{E})$ is a zero cycle supported on a curve $C\in|L|$. Since the genus of $C$ is less than or equal to $\frac{1}{2}(L^{2}+2)=\mathrm{d}(\mathcal{E})$, the assertion follows from [22, Claim 0.2] (see also [28, Proposition 15]). ∎
@@
Due to Proposition 3.3, $\pi_{1}$ is dominant. By the same argument in [22, Proposition 1.3], we know that $\pi_{2}$ is also dominant.
@@
Then for any $\mathcal{F}\in\pi_{2}(\pi_{1}^{-1}(\Sigma_{\mathscr{X}}^{\prime}))$, we have $c_{2}(\mathcal{F})\in\mathbf{S}_{0}(X)$ by [28, Theorem 9]. Therefore $c_{2}(\mathcal{E}^{\prime})\in\mathbf{S}_{0}(X)$ by the definition of $\Gamma$ for any $\mathcal{E}^{\prime}\in\Sigma_{\mathscr{X}}^{\prime}$. ∎
@@ 3.3. Correspondence between torsion sheaves
If $\mathscr{X}\to X$ is essentially trivial, this is established in [17, Theorem 3.1 & Theorem 1.1]. If $\mathscr{X}\to X$ is not essentially trivial, we can establish a finite correspondence between $M_{X}^{\operatorname{sm}}$ and $M_{\mathscr{X}}^{\operatorname{sm}}$ which preserves the two filtrations. More precisely, consider the incidence variety
@@
with $\dim Z\geq\mathrm{d}(\mathcal{E})-i$. Then $Z^{\prime}=\pi_{2}(\pi_{1}^{-1}(Z))$ is a constant cycle subvariety by Theorem 2.3 and we have $\dim Z^{\prime}=\dim Z\geq\frac{\mathbf{v}^{2}+2}{2}-i$. This means $\dim O_{E}\geq\frac{\mathbf{v}^{2}+2}{2}-i$ for any $E\in Z^{\prime}$. According to the proof of [17, Theorem 1.1], we have
@@
Let us explain the construction of $W$. Let $\Sigma_{\mathscr{X}}\subseteq M_{\mathscr{X}}$ and $\Sigma_{X}\subseteq M_{X}$ be constant cycle Lagrangian subvarieties defined in the proof of Proposition 3.1. Their restrictions to the smooth locus $M_{\mathscr{X}}^{\operatorname{sm}}$ (resp. $M_{X}^{\operatorname{sm}}$) are multisections of the abelian fibration of degree $d_{1}$ (resp. $d_{2}$).
@@ 4.1. Construction of $\mathfrak{o}_{\mathscr{X}}$
$\mathfrak{o}_{\mathscr{X}}=\frac{c_{2}(\mathcal{F})}{\operatorname{rank}\mathcal{F}}+(1-\frac{\deg(c_{2}(\mathcal{F}))}{\operatorname{rank}\mathcal{F}})\mathfrak{o}_{X},$ (18)
which is a zero cycle of degree $1$.
@@
Let $\mathbf{v}\in\widetilde{\mathrm{H}}(\mathscr{X})$ and assume $\mathbf{v}^{\perp}\subseteq\widetilde{\mathrm{H}}_{\rm alg}(\mathscr{X})$ contains an isotropic vector. Assume that $\operatorname{rank}(\mathbf{v})=r>\frac{\mathbf{v}^{2}+2}{2}$ and there exists a slope stable element in $M_{H}(\mathscr{X},\mathbf{v})$. Then $M_{H}(\mathscr{X},\mathbf{v})$ contains a constant cycle Lagrangian subvariety $Z$ such that $Z\cap M_{H}^{\rm l.f.}(\mathscr{X},\mathbf{v})\neq\emptyset$.
@@
Note that under the condition $r>\frac{\mathbf{v}^{2}+2}{2}$, all $\mu_{H}$-stable $\mathscr{X}$-twisted sheaves are locally free (cf[29, Remark 3.2] and [11, Remark 6.1.9]) and the assertion follows from our assumption.
Slope stable elements can be readily obtained. Let $r_{0}$ be the minimal rank of $\mathscr{X}$-twisted sheaves in $\mathrm{Coh}^{(1)}(\mathscr{X})$. Then $r_{0}$ equals the order of $[\mathscr{X}]$ in $\operatorname{Br}(X)$, and every element in $\mathrm{Coh}^{(1)}(\mathscr{X})$ has rank $ar_{0}$ for some integer $a\in\mathbb{Z}$ (cf. [30, Lemma 3.2]). For any primitive vector
$\mathbf{v}=(ar_{0},h,s)\in\widetilde{\mathrm{H}}(\mathscr{X})$
with $a>0$ and $\gcd(a,H\cdot h)=1$, the sheaves $\mathcal{E}\in M_{H}(\mathscr{X},\mathbf{v})$ are $\mu_{H}$-stable. When $\mathrm{rank}(\mathbf{v})>\frac{\mathbf{v}^{2}+2}{2}$, we know that every element in $M_{H}(\mathscr{X},\mathbf{v})$ is locally free.
To ensure the existence of $\mathbf{v}$ with isotropic vector in its orthogonal complement, set $s=0$. Then $(1,0,0)\in\mathbf{v}^{\perp}$ is isotropic.
@@
There is a more natural perspective on the construction of $\mathfrak{o}_{\mathscr{X}}$: it can be obtained via the second Chern class of the Beauville-Voisin class on a Fourier-Mukai partner. This perspective applies specifically when $N=M_{H}(\mathscr{X},\mathbf{w})$ is a K3 surface, in which case the local freeness requirement for $\mathcal{F}$ (lying on a constant cycle Lagrangian subvariety) becomes unnecessary.
Let $\mathcal{F}\in N$ be any point on a constant cycle curve $C\subseteq N$. In this case, $\mathbf{S}_{0}^{\mathrm{BV}}\operatorname{CH}_{0}(N)$ is known to be $1$-dimensional and hence equal to $\mathbf{S}_{0}^{\mathrm{SYZ}}\operatorname{CH}_{0}(N)$. By Theorem 1.1, for any $\mathcal{F}\in C\subset N$, we have:
$c_{2}(\mathcal{F})-\operatorname{rank}(\mathcal{F})\mathfrak{o}_{\mathscr{X}}=(\deg(c_{2}(\mathcal{F}))-\operatorname{rank}(\mathcal{F}))\mathfrak{o}_{X}.$
This implies:
$\mathfrak{o}_{\mathscr{X}}=\frac{c_{2}(\mathcal{F})}{\operatorname{rank}\mathcal{F}}+\left(1-\frac{\deg(c_{2}(\mathcal{F}))}{\operatorname{rank}\mathcal{F}}\right)\mathfrak{o}_{X}$ (19)
provided $\mathrm{rank}(\mathbf{w})>0$. For $\dim N>2$, it remains open whether the equation (19) holds for arbitrary $\mathcal{F}$ lying on a constant cycle Lagrangian subvariety.
Remark 4.5 (Descent property).
We speculate that $\mathfrak{o}_{\mathscr{X}}=\mathfrak{o}_{X}$. The main evidence is that due to our construction, the twisted BV class $\mathfrak{o}_{\mathscr{X}}$ has nice descent property. This follows from the descent property of constant cycle curves on K3 surfaces (cf[8, Proposition 3.6]).
To be precise, suppose $X$ is defined over $\bar{k}$. Take a Mukai vector $\mathbf{v}=(r,h,s)$ such that $\mathbf{v}^{2}=0$ and $r>0$. Then $M=M_{H}(\mathscr{X},\mathbf{v})$ is a K3 surface defined over $\bar{k}$ and it consists of Gieseker stable $\mathscr{X}$-twisted locally free sheaves. Take a $\bar{k}$-rational curve $C\subseteq M$ and let $\mathcal{F}$ be a point on $C$. Then $\mathfrak{o}_{\mathscr{X}}=\frac{c_{2}(\mathcal{F})}{r}+(1-\frac{\deg(c_{2}(\mathcal{F}))}{r})\mathfrak{o}_{X}$ is clearly defined over $\bar{k}$.
When $\bar{k}=\bar{\mathbb{Q}}$, the Bloch-Beilinson conjecture predicts $\operatorname{CH}_{0}(X_{\bar{\mathbb{Q}}})=\mathbb{Z}\mathfrak{o}_{X}$ and hence $\mathfrak{o}_{\mathscr{X}}=\mathfrak{o}_{X}$.
In the rest, we will show that $\mathfrak{o}_{\mathscr{X}}$ is independent of the choice of the moduli space $N$ and the constant cycle Lagrangian subvariety $Z$. The key result is
@@
The idea is to consider the degeneracy loci of morphisms between twisted sheaves. Let us recall the definition and some basic properties.
@@
Let $\mathcal{E}$ and $\mathcal{F}$ be two $\mathscr{X}$-twisted locally free sheaves. For $t:\mathcal{E}\to\mathcal{F}\in\operatorname{Hom}(\mathcal{E},\mathcal{F})$, we define the $k$-th degeneracy locus $\mathscr{D}_{k}(t)$ to be the closed substack of $\mathscr{X}$ defined by the image of the map
@@
We can also define the schematic $k$-th degeneracy locus $D_{k}(t)$ which is the scheme-theoretic image in $X$. (cf[18, Definition 2.2.6.2])
Form the definition, since $\mathscr{H}om(\wedge^{k+1}\mathcal{E},\wedge^{k+1}\mathcal{F})$ is the pullback of a sheaf on $X$, $\mathscr{D}_{k}(t)\to D_{k}(t)$ is the pullback of $\mathscr{X}\to X$ via $D_{k}(t)\to X$. The following result is well-known for untwisted locally free sheaves (cf[7, §4.1][14, Lemma 2.14]).